OpenStudy (pphalke):

The Jurassic Zoo charges $11 for each adult admission and $2 for each child. The total bill for the 233 people from a school trip was $1069. How many adults and how many children went to the zoo?

OpenStudy (pphalke):

so far i did x+y=233 then 11x+2y=1069

OpenStudy (pphalke):

what do i do next?

OpenStudy (pphalke):

@SerenityJustice74 @ace-n-it @Scooby6363

OpenStudy (pphalke):

@HyperPiper

OpenStudy (pphalke):

@ganeshie8

jimthompson5910 (jim_thompson5910):

you have the correct equations set up x = # of adults y = # of children

jimthompson5910 (jim_thompson5910):

here's what I'd do first solve for y in x+y=233 x+y=233 x+y-x=233-x y = 233 - x I subtracted x from both sides with me so far?

OpenStudy (pphalke):

yes i am

jimthompson5910 (jim_thompson5910):

ok then we will plug that into the other equation

OpenStudy (pphalke):

so what would it be

jimthompson5910 (jim_thompson5910):

11x+2y=1069 11x+2( y )=1069 11x+2( 233 - x )=1069 ... notice how I replaced `y` with `233 - x` do you see how the substitution got rid of all the y terms? and how we have nothing but x left?

jimthompson5910 (jim_thompson5910):

this replacement is possible because y = 233 - x. Basically saying y and 233-x are the same thing

OpenStudy (pphalke):

what do i do from there?

jimthompson5910 (jim_thompson5910):

next we distribute the 2 through 11x+2( 233 - x )=1069 11x+2( 233 ) + 2(-x )=1069 11x + 466 - 2x = 1069 what comes next?

OpenStudy (pphalke):

9x+466=1069?

jimthompson5910 (jim_thompson5910):

yep, then what?

OpenStudy (pphalke):

9x=603

jimthompson5910 (jim_thompson5910):

good, then there's one more step to solve for x

OpenStudy (pphalke):

yea divide by 9 and it becomes x=67

jimthompson5910 (jim_thompson5910):

correct x = 67

jimthompson5910 (jim_thompson5910):

then you'll use this to find y x+y=233 67+y=233 y = ????

OpenStudy (pphalke):

you subtract 67 frm both sides right??

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (pphalke):

y=166

jimthompson5910 (jim_thompson5910):

so there are 67 adults and 166 children

jimthompson5910 (jim_thompson5910):

check: x+y = 233 67+166 = 233 233 = 233 ... works 11*x + 2*y = 1069 11*67 + 2*166 = 1069 737 + 332 = 1069 1069 = 1069 ... works both equations are true when (x,y) = (67, 166) so the solution is confirmed